# Nonlinear Schrödinger equation in terms of elliptic and hyperelliptic $\sigma$ functions

### 10 Mar 2024 v1

@@ 5. Baker function for $X$
Here we let ${{\widetilde{\zeta}}_{i}(P):=-\int^{P}_{\infty}{\nu^{\mathrm{II}}_{i}}}$ for $u\in{\mathbb{C}}^{g}$ and $P\in X$, we let $v=w(P)$.
Proposition 4.8, Lemmas 4.11 and 4.13 show the generalization of (2.11).
@@
Proof.
The same as the case of $g=1$.
@@ 6.1. Genus two
Let us consider only the genus $g=2$ case as follows. We assume $(P_{1},P_{2},{\widetilde{P}})\in S^{2}X\times X$, $(P_{i}=(x_{i},y_{i})$, and ${\widetilde{P}}=({\widetilde{x}},{\widetilde{y}})$ such that $u={\widetilde{w}}(\iota_{X}(P_{1},P_{2}))$ and $v={\widetilde{w}}(\iota_{X}({\widetilde{P}}))$.
@@
We immediately obtain the following relation from (6.1), though we don’t use it in this paper,
@@
Now we state our main theorem in this paper, in which we obtain the generalization of (2.14) or (1.1).
@@
For $(P_{1},P_{2},{\widetilde{P}})\in S^{2}X\times X$ such that 1) $x_{1}x_{2}=-{\widetilde{x}}^{2}$, or 2) $x_{1}x_{2}=0$, and $\alpha\in{\mathbb{C}}^{g}$ and $c,\beta\in{\mathbb{C}}$, we let
$\Psi(u;v,t):=\frac{1}{\sqrt{{\widetilde{x}}}}{\mathrm{e}}^{{\,{}^{t}\negthinspace}\alpha u+{\mathfrak{i}}\beta t+c}\psi(u;v),\quad\Psi^{*}(u;v,t):=\frac{-1}{\sqrt{{\widetilde{x}}}}{\mathrm{e}}^{-{\,{}^{t}\negthinspace}\alpha u-{\mathfrak{i}}\beta t-c}\psi(u;-v).$
The following formulae hold.
$\begin{split}\mbox{1)\ :\ }&\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi-(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+{\widetilde{x}}+\lambda_{2}-\alpha^{2})\Psi,\\ \mbox{2)\ :\ }&\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi-(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta-{\widetilde{x}}+\lambda_{2}-\alpha^{2})\Psi.\\ \end{split}$
Proof.
From (5.2), we have
$\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi-\wp_{gg}\Psi=\frac{1}{2}(2\beta+{\widetilde{x}}+\lambda_{2}-\alpha^{2})\Psi.$
Under the restriction 1), $(\Psi^{*}\Psi)$ is reduced to $(1/{\widetilde{x}})$ times ${\widetilde{x}}\wp_{22}(u)$. whereas under the restriction 2), $(\Psi^{*}\Psi)$ is equal to ${\widetilde{x}}-\wp_{22}(u)$.
Remark 6.3.
As mentioned in Introduction the Baker equation (5.2) contains the potential $\wp_{gg}$, while the terms Weierstrass-$\wp$ in $F({\widetilde{x}})$ in the addition theorem (4.2) do not coincide with $\wp_{gg}$ except for $g=1$. Thus, the discrepancy prevents us from obtaining the generalization of (1.1) directly for a long time.
Even for genus two, we find it by restriction the configurations of points in $S^{2}X$. In other words, the restriction may be effective for the hyperelliptic solution of the NLS equation in terms of $\sigma$ function and enable us to find the generalization to general genus.
However we should consider how the restriction effects on the Jacobian $J_{X}$. Further we note that in the NLS equation, we have crucial problems related to the reality conditions depending on the ”nonfocusing” and ”focusing” NLS equations [P] or the choice of $(\alpha,u)$ which belongs to ${\mathfrak{i}}{\mathbb{R}}\times{\mathbb{R}}$ or ×${\mathbb{R}}\times{\mathfrak{i}}{\mathbb{R}}$.. However, since these problems are basically complicated as in [Ma4], we will consider these problems as the next steps. we have considered the similar problem for the MKdV equation,
@@ 6.2. Genus three
Let us consider the genus $g=3$ case as follows. We assume $(P_{1},P_{2},P_{3}{\widetilde{P}})\in S^{3}X\times X$, $(P_{i}=(x_{i},y_{i})$, and ${\widetilde{P}}=({\widetilde{x}},{\widetilde{y}})$ such that $u={\widetilde{w}}(\iota_{X}(P_{1},P_{2},P_{3}))$ and $v={\widetilde{w}}(\iota_{X}({\widetilde{P}}))$.
@@
Now we state our main theorem in this paper, in which we obtain the generalization of (2.14) or (1.1).
@@
$\Psi(u;v,t):=\frac{1}{{\widetilde{x}}}{\mathrm{e}}^{{\,{}^{t}\negthinspace}\alpha u+{\mathfrak{i}}\beta t+c}\psi(u;v),\quad\Psi^{*}(u;v,t):=\frac{-1}{{\widetilde{x}}}{\mathrm{e}}^{-{\,{}^{t}\negthinspace}\alpha u-{\mathfrak{i}}\beta t-c}\psi(u;-v).$
The following formula holds.
$\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi-(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta-{\widetilde{x}}+\lambda_{2}-\alpha^{2})\Psi,$
where $\wp_{33}={\widetilde{x}}(\gamma_{1}+\gamma_{2}+1/\gamma_{1}\gamma_{2})$.
Proof.
From (5.2), we have
$\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi-\wp_{gg}\Psi=\frac{1}{2}(2\beta+{\widetilde{x}}+\lambda_{2}-\alpha^{2})\Psi.$
Under the restriction, $(\Psi^{*}\Psi)$ is reduced to $(1/{\widetilde{x}})^{2}$ times ${\widetilde{x}}^{3}-{\widetilde{x}}^{2}\wp_{33}(u)$.
@@ Acknowledgments:
The author is grateful to Professor Emma Previato for suggestions on this study at a very early stage and to Dr. Kayo Kinjo for discussing the algebraic solution of the NLS equation based on [BBEIM]. The author has been supported by the Grant-in-Aid for Scientific Research (C) of Japan Society for the Promotion of Science Grant, No.21K03289.
Competing interests:
The author declares none.

### 14 Mar 2024 v2

@@ 5. Baker function for $X$
Here ${{\widetilde{\zeta}}_{i}(P):=-\int^{P}_{\infty}{\nu^{\mathrm{II}}_{i}}}$ for $u\in{\mathbb{C}}^{g}$ and $P\in X$, and $v={\widetilde{w}}(\iota_{X}P)$. We define it as a meromorphic function of $J_{X}\times W_{X}^{1}$ while Buchstaber, Enolskii, and Leykin defined it as $J_{X}\times J_{X}$ [BEL1]. Thus ours is different from the one in [BEL1].
Proposition 4.5, Lemmas 4.8 and 4.10 show the generalization of (2.11).
@@
Proof:  The same as the case of $g=1$.
@@ 6.1. Genus two
Let us consider only the genus $g=2$ case as follows: We assume $(P_{1},P_{2},{\widetilde{P}})\in S^{2}X\times X$, $(P_{i}=(x_{i},y_{i})$, and ${\widetilde{P}}=({\widetilde{x}},{\widetilde{y}})$ such that $u={\widetilde{w}}(\iota_{X}(P_{1},P_{2}))$ and $v={\widetilde{w}}(\iota_{X}({\widetilde{P}}))$.
@@
We immediately obtain the following relation from (6.1), though we do not use it in this paper,
@@
Now we state our first main theorem in this paper, in which we obtain the generalization of (2.14) or (1.1).
@@
For $(P_{1},P_{2},{\widetilde{P}})\in S^{2}X\times X$ such that 1) $x_{1}x_{2}=-{\widetilde{x}}^{2}$, or 2) $x_{1}x_{2}=0$, and $\alpha\in{\mathbb{C}}^{2}$ and $c,\beta\in{\mathbb{C}}$, we let
$\Psi(u;v,t):=\frac{1}{\sqrt{{\widetilde{x}}}}{\mathrm{e}}^{{\,{}^{t}\negthinspace}\alpha u+{\mathfrak{i}}\beta t+c}\psi(u;v),\quad\Psi^{*}(u;v,t):=\frac{1}{\sqrt{{\widetilde{x}}}}{\mathrm{e}}^{-{\,{}^{t}\negthinspace}\alpha u-{\mathfrak{i}}\beta t-c}\psi(u;-v).$
The following formulae $(g=2)$ hold for the conditions 1) and 2),
1) : $\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi-\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+\alpha_{g}^{2}-{\widetilde{x}}-\lambda_{2})\Psi,$
2) : $\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi-\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+\alpha_{g}^{2}-3{\widetilde{x}}-\lambda_{2})\Psi.$
Proof:  From (5.2), we have
$\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi-\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi+\wp_{gg}\Psi=\frac{1}{2}(2\beta+\alpha_{g}^{2}-{\widetilde{x}}-\lambda_{2})\Psi.$
Under the restriction 1), $(\Psi^{*}\Psi)$ is reduced to $(1/{\widetilde{x}})$ times ${\widetilde{x}}\wp_{22}(u)$. whereas under the restriction 2), $(\Psi^{*}\Psi)$ is equal to $-({\widetilde{x}}-\wp_{22}(u))$.
@@ 6.2. Genus three
Let us consider the genus $g=3$ case as follows. We assume $(P_{1},P_{2},P_{3},{\widetilde{P}})\in S^{3}X\times X$, $(P_{i}=(x_{i},y_{i})$, and ${\widetilde{P}}=({\widetilde{x}},{\widetilde{y}})$ such that $u={\widetilde{w}}(\iota_{X}(P_{1},P_{2},P_{3}))$ and $v={\widetilde{w}}(\iota_{X}({\widetilde{P}}))$.
@@
Now we state our second main theorem in this paper, in which we obtain the generalization of (2.14) or (1.1).
@@
$\Psi(u;v,t):=\frac{1}{{\widetilde{x}}}{\mathrm{e}}^{{\,{}^{t}\negthinspace}\alpha u+{\mathfrak{i}}\beta t+c}\psi(u;v),\quad\Psi^{*}(u;v,t):=\frac{1}{{\widetilde{x}}}{\mathrm{e}}^{-{\,{}^{t}\negthinspace}\alpha u-{\mathfrak{i}}\beta t-c}\psi(u;-v).$
The following formula $(g=3)$ holds.
$\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi-\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+\alpha_{g}^{2}-3{\widetilde{x}}-\lambda_{4})\Psi.\$
Proof:  From (5.2), we have
$\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha_{g}\frac{\partial}{\partial u_{g}}\right]\Psi-\frac{1}{2}\frac{\partial^{2}}{\partial u_{g}^{2}}\Psi+\wp_{gg}\Psi=\frac{1}{2}(2\beta+\alpha_{g}^{2}-{\widetilde{x}}-\lambda_{4})\Psi.$
Under the restriction, $(\Psi^{*}\Psi)$ is reduced to $(1/{\widetilde{x}})^{2}$ times $-({\widetilde{x}}^{3}-{\widetilde{x}}^{2}\wp_{33}(u))$.
Remark 6.4.
As mentioned in Introduction the Baker equation (5.2) contains the potential $\wp_{gg}$, while the terms Weierstrass-$\wp$ in $F({\widetilde{x}})$ in the addition theorem (4.3) do not coincide with $\wp_{gg}$ except for $g=1$. Thus, this discrepancy prevents us from obtaining the generalization of (1.1) directly for a long time.
Even for genera two and three, we found the generalization by restricting the configurations of points in $S^{g}X$. In other words, the restriction may be effective for the hyperelliptic solution of the NLS equation in terms of $\sigma$ function, allowing us to easily find a generalization to the general genus.
However we should consider how the restriction effects on the Jacobian $J_{X}$. Further we note that in the NLS equation, we have crucial problems related to the reality conditions depending on the “nonfocusing” and “focusing” NLS equations [P] or the choice of $(\alpha,u)$ which belongs to ${\mathfrak{i}}{\mathbb{R}}\times{\mathbb{R}}$ or ×${\mathbb{R}}\times{\mathfrak{i}}{\mathbb{R}}$. Since these problems are basically complicated as in [Ma4], we will consider them as the next steps.
@@ Acknowledgments:
The author is grateful to Professor Emma Previato for suggestions on this study at a very early stage and to Dr. Kayo Kinjo for discussing the algebraic solution of the NLS equation based on [BBEIM]. The author has been supported by the Grant-in-Aid for Scientific Research (C) of the Japan Society for the Promotion of Science Grant, No.21K03289.